OpenStudy (anonymous):

solve! 12/n-1=10/n+3 n= -18 n= -23 n= -30 or n= -46

Parth (parthkohli):

12(n + 3) = 10(n - 1) Solve now.

OpenStudy (anonymous):

I don't know how...

Parth (parthkohli):

http://www.studygs.net/equations.htm Here is a tutorial.

OpenStudy (anonymous):

-23

OpenStudy (anonymous):

Use parenthesis btw 12/n-1=10/n+3 is not the same as 12/(n-1)=10/(n+3) Also if you take a test are you allowed to use a graphing calculator?

OpenStudy (anonymous):

12/(n-1)=10/(n+3) u can cross multiply which is 12n+36=10n-10 that how he got -23